OpenStudy (anonymous):

Given the equation y − 3 = one half(x + 6) in point-slope form, identify the equation of the same line in standard form. x − 2y = −12 y = one halfx y = one halfx + 6 y = one halfx + 9

OpenStudy (bahrom7893):

y-3=(1/2)x+3 y=(1/2)x+6

OpenStudy (experimentx):

the first one .... also the above is also correct

OpenStudy (bahrom7893):

oh wait it wasn't standard

OpenStudy (bahrom7893):

it's the first one

OpenStudy (anonymous):

@experimentX How do you get that?